FAR-INFRARED LASER STARK SPECTROSCOPY OF METHANOL ISOTOPOMERS

\underline{M. JACKSON}, G. R. SUDHAKARAN, Department of Physics, University of Wisconsin, La Crosse, WI 54601; A. SILVEIRA Jr., Department of Chemistry, SUNY at Oswego, Oswego, NY 13126; I. MUKHOPADHYAY, Laser Programme, Centre for Advanced Technology, Indore, India, 452 013; and R. M. LEES, Department of Physics, University of New Brunswick, Fredericton, N.B., Canada E3B 5A3.

The Stark spectra of several methanol isotopomers (CH_{3}OD, ^{13}CD_{3}OH, CD_{3}OH, and CD_{3}OD) have been investigated in the far-infrared region using the HCN and DCN lasers. The spectra were recorded at room temperature for both parallel and perpendicular polarizations using electric fields up to 60 000 V/cm. Numerous transitions have been observed with these laser lines for all isotopomers. The following transitions will be presented: for CH_{3}OD, J_{K} = 6_{4} ← 5_{3} E_{2} υ_{t} = 0, J_{K} = 18_{1} ← 18_{0} E_{2} υ_{t} = 1, J_{K} = 14_{6} ← 13_{5} A υ_{t} = 1; for ^{13}CD_{3}OH, J_{K} = 18_{2} ← 17_{1} A^{−} υ_{t} = 0; and for CD_{3}OH, J_{K} = 14_{1} ← 15_{0} E_{2} υ_{t} = 1.

Analysis of the recently recorded FTFIR spectra for CD_{3}OD has yielded accurate zero-field frequencies belonging to the Q-branch multiplet for K = 11 ← 10 E_{2} υ_{t} = 0. By combining these results with the previously reported Stark fields for the J = 11 through 19 members, we have determined the electric dipole moment of this important molecular species for the first time.